# Resonant Solitary States in Complex Networks

### v1

@@ Abstract
Partially synchronized solitary states occur frequently when a synchronous system of networked oscillators is perturbed locally. Remarkably, several asymptotic states of different frequencies can coexist at the same node. Here, we uncover the mechanism underlying this multistability: The resonant back-reaction of the network’s modes on the solitary can lead to a large energy transfer between them, and the solitary’s frequency adaptation can tune the system to this resonance. We provide a full analytic analysis of this mechanism, and show which network structures enable its presence.
@@
Oscillator systems are ubiquitous in the description of natural and human-made systems. Important extended systems as diverse as the heart, the brain, firefly populations, chemical reactions, and power grids, can be modelled as networks of interconnected oscillators Strogatz (2000); Acebrón et al. (2000, 2005); Pecora et al. (2014); Rodrigues et al. (2016). One of the most fundamental common aspects of the collective behavior of such systems is synchronization Pikovsky et al. (2001). In some, such as the power grid, it is essential for their proper functioning, in others, such as the brain, it can indicate severe dysfunction. The tendency to synchronize, and thus the function of the system, is strongly influenced by the underlying network’s topology.
@@
The paradigmatic models for understanding this relationship between structure and function in coupled oscillator networks are the Kuramoto model Kuramoto (1984) and its variants. These feature extremely rich collective behavior, such as chimera states, frequency clusters, isolated desynchronization, and spatial chaos Strogatz (2000); Acebrón et al. (2000, 2005); Pecora et al. (2014); Rodrigues et al. (2016). The Kuramoto model with inertia Tanaka et al. (1997a, b) has been developed independently to study synchronization properties in power grids Bergen and Hill (1981); Witthaut et al. (2022) and biological systems Ermentrout (1991).
@@
The transition from decoupling to synchrony in inertial Kuramoto networks is characterized by spatial chaos, a form of extreme multistability Omelchenko et al. (2011). This regime typically has several states with similarly sized basins of attraction Gelbrecht et al. (2020); Halekotte et al. (2021). Among the most prominent are frequency clusters Berner et al. (2021); Olmi et al. (2014), the simplest of which are solitary states. In a solitary state, only one or a few oscillators start to rotate at their own frequencies, while the rest of the network remains synchronized Maistrenko et al. (2014); Jaros et al. (2015). Solitary states are especially prominent for the important situation of localized large perturbations Menck et al. (2014); Hellmann et al. (2020); Jaros et al. (2018), and they are particularly likely for perturbations at leaf nodes Menck et al. (2014); Nitzbon et al. (2017). Localized perturbations are important in systems like power grids, where single component failures are common and can lead to desynchronization and blackouts UCTE .
@@
Interestingly, numerical studies have revealed that solitary nodes can exist not only at the natural frequency of the oscillators, but also at shifted frequencies in networks with losses Hellmann et al. (2020) and at intermediate frequencies Nitzbon et al. (2017). In addition, the presence of multistability in these solitary nodes has been identified, allowing for the coexistence of both intermediate and natural solitary states Nitzbon et al. (2017); Halekotte et al. (2021) at the same node.
@@
Previous work has discussed the role of a networks’ topology in understanding its overall response to perturbations and the emergence of various stable states Rohden et al. (2012); Menck et al. (2014); Rodrigues et al. (2016); Nitzbon et al. (2017); Zhang et al. (2019). In some contexts, the frequency and stability of the solitary states have been studied in terms of decoupling and entrainment arguments Menck (2014); Hellmann et al. (2020). These explanations do not account for either intermediate frequency solitaries or the coexistence of several solitary states at the same node. In the following, we use averaging theory Guckenheimer and Holmes (2002) and linear response to develop a theory of the resonant coupling between the synchronized cluster and the solitary. We uncover that the described phenomena arise from the cluster’s resonantly excited complex network structure. This leads to a non-zero mean energy flow between cluster and solitary, and a frequency shift (Fig. 1). While works like Gao and Efstathiou (2018); Yue et al. (2020); Munyayev et al. (2022) have previously considered the interaction between individual oscillators and the mean field, such an interaction between collective network modes and individual oscillators has not been previously described.
@@
Figure 1: Frequencies ($\dot{\varphi}_{i}$) and time-averaged energy flow $p_{s}$ between solitary and synchronous component for a complex oscillator network following two single-node frequency perturbations, see supplementary material (SM) for an animation and the network topology (Fig. S2) Sup . The first perturbation brings the system from the synchronized regime into an intermediate solitary state in resonance with the synchronized cluster. The second perturbation causes the transition to a natural solitary state, where the solitary node is effectively decoupled and the oscillations of the synchronized cluster are much smaller. The distinct blue trajectory belongs to the solitary node 75 in Fig. 2.
@@
In the following, we will use the Kuramoto model with inertia. This is given by the coupled second-order equations
@@
where $\varphi_{i}$ are the phases of the $N$ oscillators, $P_{i}$ the driving powers, the $m_{i}$ are inertia constants, and the $\alpha_{i}>0$ are damping coefficients. The coupling is given in terms of the weighted coupling matrix $\left\{\kappa_{ij}\right\}$ as $p_{ij}(\varphi_{i},\varphi_{j})=\kappa_{ij}\sin(\varphi_{i}-\varphi_{j})$.
@@
In a synchronous state, $\varphi_{i}^{*}$, $\dot{\varphi}_{i}^{*}=\hat{\omega}$, the phases satisfy $P_{i}-\alpha_{i}\hat{\omega}=\sum_{j=1}^{N}p_{ij}(\varphi_{i}^{*},\varphi_{j}^{*})$. By going to a corotating frame, we can always set $\hat{\omega}=0$, and will assume so from now on. The $p_{ij}(\varphi_{i}^{*},\varphi_{j}^{*})$ can be interpreted as the energy flowing through the network to balance out the driving powers $P_{i}$, (this is the physical interpretation for power grids), and $\alpha_{i}$ determines the restoring force of the frequency. The natural frequencies of the uncoupled oscillators (i.e., when all $\kappa_{ij}=0$) are determined by the condition that driving power and restoring force balance: $\Omega_{i}\coloneqq P_{i}/\alpha_{i}$.
Figure 2 shows the asymptotic frequencies actually observed following single node perturbations in a complex network generated with Schultz et al. (2014) for which $\Omega_{i}=\pm 10$. When perturbations occur at nodes 74, 75 and 84, they induce intermediate frequencies. These nodes are leaves with a highly connected neighbor. Further, intermediate frequencies occur for all such nodes in this network.
@@
Figure 2: Asymptotic solitary frequencies in a synthetic power grid after perturbations. Three solitary states are clearly distinct from the natural frequencies at ±$\pm 10$. They occur at leaf nodes with highly connected neighbors (cf. Nitzbon et al. (2017)) nodes 74, 75 and 84, shown in the insets. Nodes shown as ’$+$’ (’--’) have $P_{i}=1$ $(-1)$. States are counted as desynchronized if there are frequencies whose absolute value exceeds the threshold $\omega_{th}=1$, and as solitary if there is exactly one desynchronized node, cf. Fig. S4. Shown here is the absolute value of $\omega_{s}$, which coincides with $\omega_{s}\cdot\text{sgn}(P_{i})$ for each node $i$. See Fig. S2 for the full network, and Fig. 1 for a solitary state of node 75.
@@
Ansatz and averaging.— We will begin by introducing an ansatz for describing a solitary leaf node $z$ and a synchronized cluster $S=\{i:1\leq i\leq N,i\neq z\}$. For simplicity, we will assume that $\alpha_{i}$ and $m_{i}$ are constant throughout $S$. Node $z$ is connected to $S$ through the intermediary node $k\in S$. The coupling between them is given by $p_{zk}$, which we decompose into a long-term time average $p_{s}=\langle p_{zk}\rangle$ and a mean zero part: $p_{zk}(t)\coloneqq p_{s}+p_{zk}^{\text{osc}}(t)$. Our ansatz is
$\varphi_{i}(t)$ $=\omega_{\text{sync}}t+\vartheta_{i}^{*}+\vartheta_{i}(t)\quad\text{for $i\in S$,}$ (2)
$\varphi_{z}(t)$ $=\omega_{s}t+\vartheta_{k}^{*}+\vartheta_{z}(t).$ (3)
Here, we define $\omega_{\text{sync}}$ and $\omega_{s}$ as the long term average frequency of the synchronized cluster, and the solitary oscillator, respectively. The $\vartheta_{i}^{*}$ are a synchronous state of $S$ given a fixed power injection $p_{s}$ at $k$.
Solitary states are those for which $\vartheta_{i}(t)$ and $\vartheta_{z}(t)$ are bounded and the corotating solitary frequency, defined as $\omega_{c}\coloneqq\omega_{s}-\omega_{\text{sync}}$, is non-zero. It can also be shown by summing over Eq. (1), that $(N-1)\alpha\omega_{\text{sync}}+\alpha_{z}\omega_{s}=0$. This gives us
Now, a central step is to linearize the equations around $\vartheta_{i}=0$. This is justified numerically, as the synchronized oscillators remain relatively close to the synchronous state. To linear order in $\vartheta_{k}$, $p_{zk}$ is
$p_{zk}(t)$ $=\kappa_{zk}\left\{\sin\left[\omega_{c}t+\vartheta_{z}(t)\right]-\vartheta_{k}(t)\cos\left[\omega_{c}t+\vartheta_{z}(t)\right]\right\}.$ (5)
Introducing the vectors $e^{k}_{i}:=\delta_{ik}$, and $\vec{\vartheta}$ with components $\vartheta_{i}$ for all $i\in S$, the system is

### v3

@@ Abstract
Partially synchronized solitary states occur frequently when a synchronized system of networked oscillators with inertia is perturbed locally. Several asymptotic states of different frequencies can coexist at the same node. Here, we reveal the mechanism behind this multistability: additional solitary frequencies arise from the coupling between network modes and the solitary oscillator’s frequency, leading to significant energy transfer. This can cause the solitary node’s frequency to resonate with a Laplacian eigenvalue. We analyze which network structures enable this resonance and explain longstanding numerical observations. Another solitary state that is known in the literature is characterized by the effective decoupling of the synchronized network and the solitary node at the natural frequency. Our framework unifies the description of solitary states near and far from resonance, allowing to predict the behavior of complex networks from their topology.
I Introduction
@@
Many natural and human-made systems are characterized by various degrees of synchronization, which is one of the most fundamental common aspects of their collective behavior Pikovsky et al. (2001). Therefore, important extended systems as diverse as the heart, the brain, firefly populations, chemical reactions, and power grids, are ubiquitously modelled as networks of interconnected oscillators Strogatz (2000); Acebrón et al. (2000, 2005); Pecora et al. (2014); Rodrigues et al. (2016). In some, such as power grids, global synchronization is essential for their proper functioning, in others, such as the brain, it can indicate severe dysfunction. The tendency to synchronize, and thus the function of the system, is strongly influenced by the underlying network’s topology.
@@
The paradigmatic models for understanding this relationship between structure and function in coupled oscillator networks are the Kuramoto model Kuramoto (1984) and its variants. These models feature extremely rich collective behavior, such as chimera states, frequency clusters, isolated desynchronization, and spatial chaos Strogatz (2000); Acebrón et al. (2000, 2005); Pecora et al. (2014); Rodrigues et al. (2016); Munyayev et al. (2023); Kovalenko et al. (2021); Kuramoto (1984). The Kuramoto model with inertia Tanaka et al. (1997a, b) has been developed independently to study synchronization properties in power grids Bergen and Hill (1981); Witthaut et al. (2022) and biological systems Ermentrout (1991).
@@
The transition from decoupling to synchrony in Kuramoto networks is characterized by spatial chaos, a form of extreme multistability Omelchenko et al. (2011). This regime typically has several states with similarly sized basins of attraction Gelbrecht et al. (2020); Halekotte et al. (2021). Among the most prominent are frequency clusters Berner et al. (2021); Olmi et al. (2014), the simplest of which are solitary states. In a solitary state, only one or a few independent oscillators are phase-shifted Maistrenko et al. (2014) or, in the presence of inertia, start to rotate at their own frequencies Jaros et al. (2015), while the rest of the network remains synchronized. Solitary states are especially prominent for the important situation of localized large perturbations Menck et al. (2014); Hellmann et al. (2020); Jaros et al. (2018), and they are particularly likely for perturbations at leaf nodes Menck et al. (2014); Nitzbon et al. (2017). Localized perturbations are important in systems like power grids, where single component failures are common and can lead to desynchronization and blackouts UCTE .
@@
Interestingly, numerical studies have revealed that solitary nodes can exist not only at the natural frequency of the oscillators, but also at intermediate frequencies Nitzbon et al. (2017); Halekotte et al. (2021). In addition, the presence of multistability in these solitary nodes has been identified, allowing for the coexistence of both intermediate and natural solitary states Nitzbon et al. (2017); Halekotte et al. (2021) at the same node.
@@
Previous work has discussed the role of a networks’ topology in understanding its overall response to perturbations and the emergence of various stable states Rohden et al. (2012); Menck et al. (2014); Rodrigues et al. (2016); Nitzbon et al. (2017); Zhang et al. (2019). In some contexts, the frequency and stability of the solitary states have been studied in terms of decoupling and entrainment arguments Menck (2014); Hellmann et al. (2020). These explanations do not account for either intermediate frequency solitaries or the coexistence of several solitary states at the same node.
In this work, we use averaging theory Guckenheimer and Holmes (2002) and linear response to develop a theory of the resonant coupling between the synchronized cluster and the solitary node. We uncover that the described phenomena arise from the cluster’s resonantly excited complex network structure. This leads to a non-zero mean energy flow between the synchronized cluster and the solitary node, and a frequency shift (Fig. 1). While works like Gao and Efstathiou (2018); Yue et al. (2020); Munyayev et al. (2022) have previously considered the interaction between individual oscillators and the mean field, such an interaction between collective network modes and individual oscillators has not been previously described. Here, we uncover how the network’s topology shapes the landscape of solitary states and illustrate our findings with a range of example systems, ranging from conceptual models to fully complex networks.
II Model
@@
Figure 1: Example trajectories of solitary states of intermediate and natural frequencies. Frequencies ($\dot{\varphi}_{i}$) and time-averaged energy flow $p_{s}$ (moving average over 10 time units) between solitary and synchronous component are shown for a complex oscillator network, see Sup for an animation and the network topology (Supplementary Fig. S2). The system is perturbed twice, with the timing indicated by vertical dashed lines. Both perturbations consist of an instantaneous step of $\Delta\dot{\varphi}_{75}=6$ at node 75’s frequency. The first perturbation brings the system from the synchronized regime into an intermediate solitary state in resonance with the synchronized cluster. The second perturbation causes the transition to a natural solitary state, where the solitary node is effectively decoupled and the oscillations of the synchronized cluster are much smaller. The distinct blue trajectory belongs to the solitary node 75 in Fig. 2.
@@
In the following, we will use the Kuramoto model with inertia on complex networks Tanaka et al. (1997a, b); Hellmann et al. (2020); Zhang et al. (2019); Menck et al. (2014); Halekotte and Feudel (2020); Manik et al. (2014). It is given by the coupled second-order equations
@@
where $\varphi_{i}$ are the phases of the $N$ oscillators, $P_{i}$ the driving powers, the $m_{i}$ are inertia constants, and the $\alpha_{i}>0$ are damping coefficients. The coupling is given in terms of the weighted coupling matrix $\left\{\kappa_{ij}\right\}$ as
$p_{ij}(\varphi_{i},\varphi_{j})=\kappa_{ij}\sin(\varphi_{i}-\varphi_{j}).$ (2)
Note that the coupling matrix is symmetric, $\kappa_{ij}=\kappa_{ji}$, hence the coupling function is antisymmetric, $p_{ij}=-p_{ji}$. We parametrize the model Eq. (1) such that the synchronous state coexists with various stable attractors of different degree of synchronization. In particular, we set $P_{i}=1$ for producers, $P_{i}=-1$ for consumers, $\alpha_{i}=0.1$, and $\kappa_{ij}=6$ for connected nodes $(i,j)$. This is the parameter regime of, for example, real-world power grids Menck et al. (2014); Nitzbon et al. (2017); Hellmann et al. (2020); Halekotte and Feudel (2020); Halekotte et al. (2021).
@@
To better understand the asymptotic behavior of the system, we recall a few properties of its attractors. In a synchronous state, we have $\varphi_{i}(t)=\varphi_{i}^{*}+\hat{\omega}t$, and constant frequencies $\dot{\varphi}_{i}^{*}=\hat{\omega}$, and the phases satisfy $P_{i}-\alpha_{i}\hat{\omega}=\sum_{j=1}^{N}p_{ij}(\varphi_{i}^{*},\varphi_{j}^{*})$. By going to a corotating frame via $\varphi_{i}\to\varphi_{i}-\hat{\omega}t$, we can always set $\hat{\omega}=0$, and will assume so from now on. The $p_{ij}(\varphi_{i}^{*},\varphi_{j}^{*})$ can be interpreted as the energy flowing through the network to balance out the driving powers $P_{i}$, (this is the physical interpretation for power grids), and $\alpha_{i}$ determines the restoring force of the frequency. The natural frequencies of the uncoupled oscillators (i.e., when all $\kappa_{ij}=0$) are determined by the condition that driving power and restoring force are in balance: $\Omega_{i}\coloneqq P_{i}/\alpha_{i}$. Solitary states in systems of coupled oscillators can be quite similar to the natural solitary states of uncoupled oscillators, which can serve as an approximation Menck et al. (2014).
[…]
To illustrate the phenomenon we describe and explain in this paper, we show simulation results in Fig. 2. It displays the different asymptotic frequencies actually observed following single node perturbations in a complex network, the topology of which was generated with Schultz et al. (2014). The $|\omega_{s}|$, that coincide with $\omega_{s}\cdot\text{sgn}(P_{z})$, are the absolute values of the mean asymptotic solitary frequencies $\omega_{s}$ of the trajectories $\dot{\varphi}_{z}(t)$ at the solitary nodes $z$. The trajectories are won from integrating Eq. (1) from many initial conditions generated by perturbing the synchronous state (see Niehues (2023) for code and details). Here, asymptotic states are counted as 1-solitary if there is exactly one node with frequency $|\omega_{s}|>1$, cf. Fig. S4. Note that many solitary frequencies are close to the natural frequencies at $\Omega_{i}=\pm 10$. We define such solitary states with $\omega_{s}\approx\Omega_{z}$ as natural. However, when perturbations occur at nodes 74, 75 and 84, they can also induce intermediate frequencies. We define solitary states with such intermediate mean frequency $|\omega_{s}|<|\Omega_{z}|$ as resonant, because (i) the fluctuation of the nodes in $S$ is more excited (see Fig. 1), and (ii) because we show that $\omega_{s}$ is in resonance with a network mode. The nodes 74, 75 and 84 share a specific topological property: they are leaf nodes (degree-1 nodes) with a high-degree neighbor. The intermediate resonant solitary states occur at all such nodes in this network (and other networks), and only at these nodes, discarding marginal effects. We explore this strong link between the topology and the asymptotic behavior of the nodes in the following sections.
@@
Figure 2: Local topological features determine the existence and relative occurrence of asymptotic solitary frequencies $\omega_{s}$ in a synthetic power grid observed after perturbations, see Niehues (2023). Three solitary states are clearly distinct from the natural frequencies at ±$\pm 10$; they occur at leaf nodes 74, 75 and 84 which have highly connected neighbors (cf. Nitzbon et al. (2017)) as shown in the insets. Nodes shown as ’$+$’ (’--’) have $P_{i}=1$ $(-1)$. See Fig. S2 for the full network, and Fig. 1 for a solitary state of node 75.
III Ansatz and self-consistency
Here, we present our main result: the framework that allows us to make predictions of solitary states and relate them to the network’s topology. The framework predicts the existence and stability of both natural and resonant solitary states. It consists of an ansatz and a self-consistent equation it has to fulfill.
An ansatz is a proposed form of a solution, often justified by prior knowledge about the system, and its sole purpose is to insert it and evaluate the result it produces. Here, the ansatz is a change of coordinates without loss of generality.
We begin by introducing suitable notation for describing a solitary leaf node $z$ and a synchronized cluster $S=\{i:1\leq i\leq N,i\neq z\}$. For simplicity, we will assume that $\alpha_{i}$ and $m_{i}$ are constant throughout $S$. Node $z$ is connected to $S$ through the intermediary node $k\in S$. With all introduced notations, the considered system has the form
$m\ddot{\varphi}_{i}$ $=P_{i}-\alpha\dot{\varphi}_{i}-\sum_{j=1}^{N}\kappa_{ij}\sin(\varphi_{i}-\varphi_{j}),\quad i\in S,$ (3)
$m_{z}\ddot{\varphi}_{z}$ $=P_{z}-\alpha_{z}\dot{\varphi}_{z}-\kappa_{zk}\sin(\varphi_{z}-\varphi_{k}).$ (4)
@@ III Ansatz and self-consistency
The coupling between subsystems $S$ and $z$ is given by $p_{zk}$, which we decompose into a long-term time average $p_{s}\coloneqq\langle p_{zk}\rangle$ and a mean zero part $p_{zk}^{\text{osc}}(t)$:
$p_{zk}(t)=\kappa_{zk}\sin\left[\varphi_{z}(t)-\varphi_{k}(t)\right]\coloneqq p_{s}+p_{zk}^{\text{osc}}(t).$ (5)
Note that we do not have an explicit expression for $p_{s}$ and $p_{zk}^{\text{osc}}(t)$ yet, but obtain them later.
Our ansatz for the trajectories, is to split the linear motion in time with the mean frequency from the mean zero part, similarly to $p_{zk}$ in Eq. (5),
$\varphi_{i}(t)$ $\coloneqq\omega_{\text{sync}}t+\vartheta_{i}^{*}+\vartheta_{i}(t)\quad\text{for $i\in S$,}$ (6)
$\varphi_{z}(t)$ $\coloneqq\omega_{\text{sync}}t+\vartheta_{k}^{*}+\vartheta_{z}(t)\coloneqq\omega_{s}t+\vartheta_{k}^{*}+\varepsilon\psi_{z}(t),$ (7)
for some small $\varepsilon$ that we determine later. We define $\omega_{\text{sync}}$ and $\omega_{s}$ as the long term average frequency of the synchronized cluster, and the solitary oscillator, respectively. This means $\langle\vartheta_{i}(t)\rangle=0=\langle\psi_{z}\rangle(t)$ by construction. The $\vartheta_{i}^{*}$ are a synchronous state of $S$ given a fixed power injection $p_{s}$ at node $k$. Note that the ansatz is without loss of generality, and can be formally considered as a change of variables. We use the form Eqs. (6-7) because if a solitary state occurs, $\psi_{z}(t)$ and $\vartheta_{i}(t)$ are bounded, and the corotating solitary frequency, defined as $\omega_{c}\coloneqq\omega_{s}-\omega_{\text{sync}}$, is non-zero, cf. Supplementary Video 1 Sup . We can use the ansatz Eqs. (5-7) to determine the mean-zero part of the solitary trajectory $\psi_{z}(t)$, however, it is more informative to determine $\omega_{s}$ and $p_{s}$ from the ansatz as a characterization of the solitary state.
To determine $\omega_{s}$, we note that a condition for the existence of a solitary state at $\omega_{s}$ is that both $\omega_{s}$ and $p_{s}$ have to fulfill a self-consistent equation. The argument goes as follows. For $\psi_{z}(t)$ to stay bounded, and for $\omega_{s}$ to be the long-term average of the solitary node’s frequency, the long-term average of $\ddot{\vartheta}_{z}=\varepsilon\ddot{\psi}_{z}$ has to be zero. Demanding this for Eq. (4) provides us with the condition $0=P_{z}-\alpha_{z}\omega_{s}-p_{s}$. We can interpret $p_{s}$ as a measure for the mean coupling of $S$ and $z$, while $\omega_{s}$ stands for their incoherence. Note that this relationship between $\omega_{s}$ and $p_{s}$ can be observed in Fig. 1: in synchrony ($\omega_{s}=0$) there is a steady high energy flow $p_{s}=P_{z}$; in the natural solitary state, the mean energy flow $p_{s}\approx 0$ and the system is effectively decoupled. For intermediate resonant solitary states, there is some intermediate amount of energy transfer $0<|p_{s}|<|P_{z}|$.
If we now can express $p_{s}$ as a function of $\omega_{s}$, the above condition for a solitary state to have well-defined frequency becomes a self-consistent equation for $\omega_{s}$:
The self-consistency function $Z(\omega_{s})$ is defined as the average change of the solitary frequency as a function of the solitary mean frequency: $Z(\omega_{s})\coloneqq\langle m_{z}\ddot{\vartheta}_{z}\rangle(\omega_{s})$. A potentially stable solitary state at $\omega_{s}$ requires the correct sign of the change in $Z$ with frequency: if a change in $\omega_{s}$ leads to a larger change in $\omega_{s}$ in the same direction, the solitary frequency is linearly unstable, and vice versa. If we imagine $\omega_{s}$ as slowly varying, we have $m_{z}\dot{\omega}_{s}=Z(\omega_{s})$. Thus, we expect the sign of the derivative $\partial_{\omega_{s}}Z(\omega_{s})$ to be informative about the stability of the solitary state, at least as a proxy. In other words, by introducing the explicitly time-dependent ansatz Eqs. (6-7), we get a non-autonomous system from Eqs. (3-4). We are looking for limit cycles of this system that correspond to fixed points in the long-term averaged dynamics. These fixed points are solutions $\omega_{s}$ of Eq. (8): mean frequencies that are allowed to persist by the dynamics, and that are stable under small perturbations.
To obtain the solutions, our strategy is as follows. We assume that we can linearize Eqs. (3-4) in $\vartheta_{i}(t)$ and $\varepsilon$ to obtain a system that quantitatively reflects the behavior of the trajectory of the solitary state. We will analyze the linearized system using averaging, and see that this is justified in many regimes. Together with the self-consistent equation, we obtain a proxy system for the description of solitary states. If the proxy system is stable, we conclude that a solitary state can (but is not guaranteed to) exist at $\omega_{s}$. If the proxy system does not admit a stable solution for specific parameters, we interpret this as evidence that a solitary state can not exist with those parameters. Numerically, we see that if the proxy system admits a stable solution, we often do find solitary states with a large basin. We interpret this to mean that the mechanism revealed by the proxy system is indeed responsible for the formation and stabilization of solitary states.
To follow the strategy outlined above and evaluate Eq. (8) to get predictions for solitary frequencies and their stability, we first need to determine $p_{s}$ as a function of $\omega_{s}$. We present a derivation in the next section, and an interpretation and alternative derivation thereafter.
IV Averaging
Here, we outline the derivation of our main result, an explicit form for $p_{s}(\omega_{s})$. For a detailed step-by-step instruction, we refer to the Supplemental Material Sup , section I (SM I).
Both $p_{s}$ and $\omega_{s}$ depend on the trajectories $\varphi_{z}(t)$ and $\varphi_{k}(t)$ of the solution to Eqs. (3-4). To obtain an explicit approximation of the solution, we will leverage a version of the averaging theorem Guckenheimer and Holmes (2002) (and references therein).
We start by giving a brief summary of the averaging method following Guckenheimer and Holmes (2002). Let $0\leq\varepsilon\ll 1$, and $x\in U\subseteq\mathbb{R}^{n}$ for a bounded set $U$. Let $f:\mathbb{R}^{n}\times\mathbb{R}\times\mathbb{R}^{+}\to\mathbb{R}^{n}$ be a $C^{r}$ function with $r\geq 2$ (at least twice continuously differentiable) that is time-periodic with period $T>0$, and that determines the non-autonomous periodic dynamical system
$\dot{x}=\varepsilon f(x,t,\varepsilon).$ (9)
Define the associated autonomous averaged system as
[…]
The averaging theorem states that a system of the form Eq. (9) can be cast into the form
$\dot{y}=\varepsilon\overline{f}(y)+\varepsilon^{2}f_{1}(y,t,\varepsilon),$ (11)
where $f_{1}$ is also of period $T$ in $t$. This can be achieved by a coordinate change. Moreover, solutions $x(t)$ of Eq. (9) and $y(t)$ of Eq. (10) that are $\varepsilon$-close stay $\varepsilon$-close on a timescale $t\sim\varepsilon^{-1}$. Further, hyperbolic fixed points of Eq. (10) have corresponding unique $\varepsilon$-close hyperbolic periodic orbits of the same stability type. We will leverage this theorem as follows: By using the ansatz Eqs. (6-7) and appropriate approximations, we cast the system Eqs. (3-4) into the form Eq. (9). In the proper coordinate frame, the system is time-periodic and slowly varying. We can then find approximate solutions by solving its corresponding averaged system for fixed points. Those solutions, in particular $\vartheta_{k}(t)$, enable us to calculate $p_{s}$ as a function of $\omega_{s}$ and close the self-consistent equation, Eq. (8).
First, we need to approximate the highly nonlinear dynamics in Eqs. (3-4) appropriately. Since the synchronized oscillators from the cluster $S$ are weakly perturbed by the solitary rotation, we assume smallness of $\vartheta_{i}(t)$, which is also justified by numerical observations, see Fig. 1 and Sup ; Niehues (2023). Therefore, we linearize the system Eqs. (3-4) with respect to $\vartheta_{i}(t)\in S$ around the origin. It is crucial not to linearize with respect to $\vartheta_{z}(t)$, because it is not bounded. For example, the coupling function $p_{zk}$ in Eqs. (4) and (5) is approximated to linear order in $\vartheta_{k}$ as
$p_{zk}(t)$ $\approx\kappa_{zk}\left\{\sin\left[\omega_{c}t+\varepsilon\psi_{z}(t)\right]-\vartheta_{k}(t)\cos\left[\omega_{c}t+\varepsilon\psi_{z}(t)\right]\right\}.$ (12)
From now on, $p_{s}$ and $p_{zk}^{\text{osc}}(t)$ represent the mean and oscillating part of the approximated $p_{zk}$. To write the obtained system in vector form, we define the vectors $e^{k}_{i}:=\delta_{ik}$, where $\delta_{ik}=1$ if $i=k$, and $\delta_{ik}=0$ else. Introducing $\vec{\vartheta}(t)$ with components $\vartheta_{i}(t)$ for all $i\in S$, the linearized system is